OpenStudy (anonymous):

Let c be any constant number. Which of the following will always be perpendicular to −4x + y = −2?

OpenStudy (anonymous):

where are the following

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y=4x+c y= -4x+c y= -1/4x +c y= 1/4x+c

OpenStudy (zzr0ck3r):

solve for y in the equation from the question and tell me what you have

OpenStudy (zzr0ck3r):

yes y=4x-2

OpenStudy (zzr0ck3r):

this has a slope of 4, and any line perpendicular will have a slope of -1/4

OpenStudy (zzr0ck3r):

so what is the answer?

OpenStudy (anonymous):

C

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

so if slope is a, then perpendicular slope is -1/a